OpenStudy (anonymous):

What are the zeros of the function f(x) = 3x^2-3x-6/3x-6? a.–1 b. 1 c. 2 d. –2

OpenStudy (anonymous):

Dear, can you please rewrite this with parenthesis to clearly indicate what you meant? 3x^2-3x-6/3x-6 \(\neq\) (3x^2-3x-6)/(3x-6) \[3x^2-3x-6/3x-6 \neq \frac{3x^2-3x-6}{3x-6}\] Then do the following: 1. Factor & simplify 2. Set equal to 0, you're looking for when y=0, when your output = 0

OpenStudy (anonymous):

\[3x^2-3x-\frac{6}{3}x-6 = 3x^2-3x-2x-6 = 3x^2-5x-6\]

OpenStudy (anonymous):

So the answer is..?

OpenStudy (anonymous):

Which did you mean?

OpenStudy (anonymous):

I think c. 2. but I`m not sure.

OpenStudy (anonymous):

No I mean which of the two equations was the original intended problem. They are most certainly NOT the same thing.

OpenStudy (anonymous):

It kinda matters, a lot

OpenStudy (anonymous):

And I did ask please :-P

OpenStudy (anonymous):

If what I think you intended and not what you literally wrote, factor out a 3 first (un-distribute a 3, if you will)

OpenStudy (anonymous):

it`s \[3x^2-3x-6 \over 3x-6\]

OpenStudy (anonymous):

Yep, see what happens when you get rid of the 3? \[\frac{x^2−x−2}{x−2}\] Think you can factor that numerator now? ;-)

OpenStudy (anonymous):

\[\frac{(x-2)(x+1)}{(x-2)}\] ( any more and I'll be giving away the answer ;-) ) Make sense?

OpenStudy (anonymous):

F.O.I.L. That numerator back if you don't believe me :)

OpenStudy (anonymous):

So the answer is -1? (: